What is the distance between the points (-3,2) and (1,0)?
1
Expert's answer
2012-09-28T07:31:20-0400
d=sqrt((1-(-3))^2+(0-2)^2)=sqrt(4^2+2^2)=sqrt(16+4)=sqrt(20)=2sqrt(5)